OpenStudy (anonymous):

Solve the problem by stating the equation you would use to solve and then solving it. Be sure to show work... In a local election, 25,800 people voted. Thius was an increase of 5% over the last election. How many people voted in thne last election? Round to the nearest whole person.

OpenStudy (anonymous):

let the qt be x 5% of x is x/100x5 = x/20 total qt. x + x/20 = 25800 21x = 25800 x 20 = 24571.428 now people can't be in fraction so 24572